Suppose z=x^2siny, x=−4s^2+t^2, y=−2st.

A. Use the chain rule to find ∂z/∂s and ∂z/∂t as functions of x, y, s and t.
∂z/∂s=

∂z/∂t=

B. Find the numerical values of ∂z/∂s and ∂z/∂t when (s,t)=(−5,−5).
∂z/∂s(−5,−5)=

∂z/∂t(−5,−5)=
1
Expert's answer
2020-10-13T19:17:02-0400

"\displaystyle z=x^2\sin y, x=−4s^2+t^2, y=−2st.\\

A. \\\textsf{By Chain rule or composite rule of differentiation,}\\

\frac{\partial z}{\partial s} = \frac{\partial z}{\partial x}\times \frac{\partial x}{\partial s}\\


\frac{\partial z}{\partial x} = 2x\sin{y}\, \frac{\partial x}{\partial s} = -8s\\

\therefore \frac{\partial z}{\partial x} = -16xs\sin{y} \\

B.\\ \begin{aligned}
\frac{\partial z}{\partial t} &= \frac{\partial z}{\partial x} \times \frac{\partial x}{\partial t} \\&= -16xs\sin{y} \times 2t = -32xst\sin{y}
\end{aligned}\\


\textsf{At}\, (s, t) = (-5, -5)\\

y = (-2)(-5)(-5) = -50\\

x = -4(-5)^2 + (-5)^2 = -100 + 25 = -75\\

\therefore \frac{\partial z}{\partial x} = -16(-75)(-5)\sin{(-50)} = 6000\sin(50) = -1574.249\,\&\\

\frac{\partial z}{\partial t} = 2(-5)\times 6000\sin(50) = -60000\sin(50)=15742.49"